OpenStudy (anonymous):

Identify all like terms: hr^2 and h^2r 2x and 3y 5x^3y^6 and 2x^6y^3 wz and zw 21m^3y^4 and -9m^3y^4

OpenStudy (anonymous):

I know that 5x^3y^6 and 2x^6y^3 wz and zw 21m^3y^4 and -9m^3y^4 Are like terms, but how about the others?

jhonyy9 (jhonyy9):

hr^2 = h*r*r h^2r = h*h*r so hope so much that you see sure what is the difference

OpenStudy (anonymous):

I don't get it.